# HyGen: Efficient LLM Serving via Elastic Online-Offline Request Co-location

### v1

@@ 4.3 SLO-Aware Offline Scheduling Policies
Offline scheduling in HyGen differs from online scheduling due to the variability of resource availability. Resources allocated to offline requests depend on the residual capacity after scheduling online requests, which fluctuates with the intensity of the online workload. Consequently, offline scheduling policies must dynamically adapt to available resources and optimize their specific objectives under such constraints. Common objectives include maximizing throughput, supporting multi-priority workloads, and, optionally, ensuring fairness Sheng et al. (2024b). The prioritization of these objectives varies depending on the chosen policy.
Prefix sharing Gim et al. (2024); Gao et al. (2024); Zheng et al. (2024a) has emerged as a widely-adopted technique to significantly improve throughput and reduce latency in LLM serving systems. The benefits are particularly pronounced in offline LLM serving, where requests can be strategically ordered to maximize prefix cache reuse Zhao et al. (2024); Zheng et al. (2024b). This optimization is especially effective for offline workloads since they have relaxed latency constraints, allowing the system to accumulate and reorder requests to maximize sharing opportunities. To showcase the extensibility and flexibility of HyGen’s design, we next introduce the SLO-aware Prefix Sharing Maximization (PSM) strategy, a throughput-optimized offline scheduling policy that leverages prefix reuse to improve resource efficiency. Recognizing the potential starvation issue inherent in naive PSM, we extend it with a fairness-enhanced variant. This improved PSM algorithm incorporates mechanisms such as freshness-aware scheduling and a unified utility function to balance throughput and fairness, mitigating starvation without compromising efficiency.
The modular architecture of HyGen enables seamless integration of advanced offline scheduling policies. In addition to PSM, the system supports other strategies, such as a default First-Come-First-Serve (FCFS) policy and a Deadline-First policy for basic scheduling needs. This flexibility allows HyGen to cater to a wide range of objectives, including multi-priority support and fairness, making it a robust and adaptable solution for hybrid serving scenarios with dynamic and heterogeneous workloads.
@@
To fully leverage prefix sharing among offline requests, HyGen employs a Prefix Sharing Maximization (PSM) strategy designed for SLO-aware offline scheduling. This strategy organizes offline requests into a prefix tree and performs Depth-First Search (DFS) to generate a sorted waiting list. HyGen then performs SLO-aware scheduling on offline requests using this sorted list in a similar manner as introduced in Section 4.1, while maintaining the original DFS order to maximize prefix sharing. This enables reducing redundant computation and improving throughput.
Despite its efficiency, the naive PSM strategy can lead to starvation for requests with minimal prefix-sharing potential. This issue can be mitigated by combining requests selected by PSM with those chosen based on freshness. The trade-off between these two selection methods can be managed using a utility ratio, ensuring a balance between efficiency and fairness. During scheduling, offline waiting requests are stored in both a prefix tree for maximum prefix sharing, and a self-balanced binary search tree (BST) sorted by freshness. When a new offline request can be scheduled from the waiting list, an offline request is selected either from the head of the queue in the DFS order of the prefix tree, or from the most stale request in the self-balanced BST, based on the utility value. The other data structure then syncs up by deleting the selected request from their lists, namely by prefix and index lookup. The formalized algorithm is detailed in Appendix A.2.
While starvation due to the strict prioritization of online requests over offline ones is unavoidable to ensure performance isolation, the enhanced PSM algorithm effectively mitigates starvation caused by prefix-sharing optimizations. These enhancements make the PSM strategy more practical for real-world deployments, retaining its efficiency while improving fairness and adaptability.
@@ 5 Performance Evaluation
Figure 4: HyGen improves serving throughput under the same SLOs.
@@ 5.1 Evaluation Setup
Implementation.
We implement HyGen on top of the open-source implementation of vLLM Kwon et al. (2023); vLLM Project (2023) and Sarathi-Serve Agrawal et al. (2024); Sarathi-Serve Project (2024), with 1,300 lines of additional code.
Testbeds.
We evaluate HyGen on two server configurations. The first server is equipped with 4 NVIDIA A100 GPUs (40GB VRAM each), while the second server contains 4 NVIDIA A40 GPUs (48GB VRAM each). Both servers are identically configured with an AMD EPYC Milan CPU (1 socket, 64 cores, SMT disabled), 256GB DDR4 RAM, and a 1.5TB NVMe SSD (PCIe 4.0).
Models.
We used Llama2-7B Touvron et al. (2023) and Qwen-14B Bai et al. (2023) models to test the performance of HyGen on varying model structures and sizes. For the two models, we used A100 and A40 GPUs, respectively.
Workloads.
For the online workload, we used the conversation trace from Azure LLM inference trace 2023 Patel et al. (2024), a one-hour trace taken from the LLM services in Azure with real-world requests and timestamps. The request rate variation is shown in Figure 1. We randomly sampled the original trace to get the desired overall QPS that suits our hardware serving capacity. Specifically, for a time duration of $T$ seconds, we would sample $T\times Q$ requests from this time window to suit a desired QPS $Q$. For the offline workload, we used arXiv summarization Cohan et al. (2018), a dataset for summarization of long documents.
@@
For interference evaluation, we focus on both Time to First Token (TTFT) and Time Between Tokens (TBT) and their respective mean and 99th percentile (P99) values. For throughput evaluation, we evaluated our system on the maximum throughput of offline requests while meeting interference SLOs.
@@
For pure online inference, we use Sarathi Agrawal et al. (2024) as our baseline. For hybrid serving, we compare HyGen with two baselines for interference and throughput evaluation, respectively: (1) Sarathi++: We implement our online-first two-phase scheduling policy on Sarathi to support hybrid serving, including the request management and preemption handling policies introduced in Section 4. (2) HyGen*: To evaluate the throughput benefit of our HyGen design, we further improve Sarathi++ to be an SLO-aware serving system, HyGen*. Besides inheriting the serving policies from Sarathi++, HyGen* serves offline requests at a specific offline QPS to control overall interference. The offline QPS is profiled using a similar design with the HyGen profiler to guarantee bounded SLO interference.
5.2 End-to-end Performance
HyGen respects latency requirements in co-locating requests. We tested HyGen on four SLO metrics, namely mean TBT, P99 TBT, mean TTFT, and mean TBT. For each metric, we set varying interference tolerance ratios as the SLO. Figure 3 shows that HyGen is able to control the interference and guarantees to meet specific SLOs across our settings. Compared with Sarathi++, an SLO-unaware system that yields the same result for all metrics and tolerance ratios, HyGen showcases efficient SLO-aware latency control.
Figure 5: Prefix sharing maximum scheduler achieves higher serving capacity.
Figure 6: SLO-aware profiler bridges the gap between batch latency and overall QoE.
@@ 5.2 End-to-end Performance
Figure 8: HyGen meets stringent SLO under various QPS settings
@@
Figure 12: Hygen meets SLOs and achieves higher throughput for CNN/DM offline trace.
Figure 13: HyGen is able to meet multiple SLOs.
Figure 14: HyGen design is robust to predictor accuracy.
Figure 15: Hygen adjusts offline throughput based on online QPS.
HyGen improves serving throughput. Figure 4 showcases the offline throughput of HyGen for various metrics and tolerance ratios. Under the same SLO, HyGen consistently achieves higher throughput compared to HyGen*, yielding up to 5.37×5.37$\times$5.37 × to 5.84×5.84$\times$5.84 × throughput gain. This verifies that our fine-grained latency predictor and SLO-aware profiler achieve higher offline serving efficiency compared to their simplified counterparts.
@@ 5.3 Performance Breakdown
We evaluate the accuracy of our latency predictor on Llama2-7B and Qwen-14B using Azure LLM Inference trace mixed with arXiv summarization dataset. Figure 7 shows the CDF of absolute error and absolute percentage error of the latency predictor. HyGen latency predictor is able to achieve a mean absolute percentage error of only 1.78% and 1.07%, verifying the high accuracy of our predictor design.
@@
To test HyGen’s compatibility with prefix sharing, we did a simulation experiment using Azure LLM Inference as the online trace and MMLU Hendrycks et al. (2021a; b) as the offline dataset, running on Llama2-7B model. In our simulation, the shared prompt prefix length between two consecutively scheduled offline requests is deducted from the original prompt length of the latter one to simulate prefix sharing. We tested the offline throughput of HyGen with and without our prefix-aware offline schedule policy. Figure 6 shows that HyGen yields up to 4×4$\times$4 × offline throughput gain with its prefix-aware scheduling policy.
@@
To demonstrate the effect of the SLO-aware profiler in HyGen, we conducted experiments where we simply set the desired mean TBT SLO as the batch latency budget. Figure 6 shows the performance gap between individual batch latencies and overall mean TBT, and that our SLO-aware profiler is able to bridge this gap and provide hybrid serving with controlled SLO.
@@
Figure 9 shows a temporal breakdown of the throughput performance of HyGen during one serving process. At runtime, HyGen shows the ability to dynamically adjust offline throughput according to current online workload and overall latency budget. HyGen batches offline requests more aggressively during troughs of online QPS and lowers offline throughput during online bursts, harnessing compute resources in an adaptive manner.
@@ 5.4 Ablation Studies
To evaluate HyGen’s effectiveness in distributed inference scenarios, we deployed the Yi-34B model Young et al. (2024) using a hybrid parallelization strategy: tensor-parallelism (TP) Shoeybi et al. (2019) and pipeline-parallelism (PP) Athlur et al. (2022); Narayanan et al. (2019); Huang et al. (2019) with degree 2 for each dimension. The online and offline workloads are Azure LLM Inference trace and arXiv summarization, respectively. Figure 10(a) and Figure 10 show that HyGen maintains its ability to meet SLOs and achieves higher throughput (up to 1.89×1.89$\times$1.89 ×) than the baseline.
@@
To further evaluate HyGen’s ability to meet stringent SLOs under various settings, we conducted experiments on the four aforementioned metrics under varied QPS settings, each with a 5% interference tolerance ratio. Figure 8 shows that HyGen is able to meet stringent SLOs for all metrics under various QPS settings.
We further demonstrate HyGen’s ability to meet multiple SLOs at the same time. Here, we tested HyGen with a fixed P99 TTFT interference ratio (8%) and mean TBT interference ratios ranging from 10% to 50%. Figure 13 shows that at a lower mean TBT tolerance, the performance of HyGen is bounded by mean TBT SLOs; however, after reaching the fixed P99 TTFT SLO, mean TBT stops increasing in order to keep P99 TTFT under control, demonstrating the flexibility of HyGen in meeting multiple SLOs.
@@
We further evaluate HyGen on Mistral-7B and MooncakeQin et al. (2024), a trace containing request lengths and timestamps taken from real-world servers, as the online trace. The offline dataset used in the first experiment is arXiv summarization. For the second experiment, we replaced the offline dataset with CNN/DailyMail summarization dataset See et al. (2017); Hermann et al. (2015) and used Llama2-7B and Azure LLM Inference trace as the model and online dataset, respectively. Figure 11 and Figure 12 show HyGen achieves superior performance than its counterparts under these two settings.
@@
To further evaluate the robustness of HyGen, we tested HyGen using several pre-trained LR latency predictors taken from other settings and workloads and tested them on Azure LLM Inference trace, resulting in varied prediction accuracy. Figure 15 shows how predictor accuracy (measured in mean absolute percentage error) affects offline throughput under the same P99 TBT SLO. We notice that HyGen remains robust across different accuracy settings. Additionally, our LR-based latency predictor is lightweight for training, with only $\sim$15ms training time for over 80,000 samples on our testing CPU, guaranteeing transferability and negligible overhead.
Impact of online arrival rates.
Under high online request arrival rates, online requests may approach system serving capacity, leaving little headroom for offline request co-location. Figure 15 shows how online QPS affects offline throughput under a 5% P99 TBT SLO tolerance ratio. As online load increases, HyGen is able to adjust offline throughput according to the residual serving capacity of the system, while maintaining higher throughputs.
@@ 6 Conclusion
This paper introduces HyGen, an LLM serving system that enables efficient co-location of online and offline workloads. We employ a control mechanism to predict and manage interference impacts, while the scheduling framework supports multiple optimization objectives for offline serving. Evaluation on production workloads demonstrates that HyGen achieves up to 5.84$\times$ higher throughput while maintaining strict latency SLOs.
@@ A.1 HyGen Two-phase Scheduling Algorithm
This section gives a detailed explanation of the two-phase scheduling algorithm introduced in Section 4. To reduce the scheduling overhead of HyGen, we employ a message queue for asynchronous communication between the main process and the offline scheduler. After each scheduling step, the main process sends the metadata of batched requests to the message queue. The offline scheduler first calculates the expected status of each request based on scheduling decisions from the previous batch, and then runs a scheduling simulation to generate offline request scheduling decisions using our latency predictor and the profiled latency budget. The offline scheduling decisions are then sent back to the main process using the message queue and used for the coming scheduling step. To support pipeline parallelization, a scheduling history archive of $K$ steps is kept by the offline scheduler for pipeline parallelization size $K$, in order to have a holistic view of every request running in each pipeline stage at the time.
@@
3:  global $Q_{send}$, $Q_{recv}$ (message queue)

### v2

@@ 4.3 SLO-Aware Offline Scheduling Policies
After scheduling online requests, our offline scheduling policy repurposes the residual capacity to maximize throughput while ensuring fairness Sheng et al. (2024b).
To further optimize the serving throughput for offline requests, HyGen employs an SLO-aware Prefix Sharing Maximization (PSM) strategy. Prefix sharing is a widely-adopted technique for reusing the KV cache of shared input prefix between requests Gim et al. (2024); Zheng et al. (2024a); Zhao et al. (2024); Zheng et al. (2024b).
@@
Our Prefix Sharing Maximization (PSM) strategy organizes offline requests into a prefix tree that follows the structure of a Trie tree, capturing prefix sharing characteristics of all offline requests, with each leaf node representing an offline request. The priority of each request is determined by the Depth-First Search (DFS) order of the prefix tree, where requests with the greatest prefix sharing potential are scheduled together, maximizing prefix cache reuse among requests. Subsequently, HyGen performs SLO-aware offline scheduling (Section 4.1) using this order to maximize prefix sharing, reducing redundant computation and improving throughput. Additionally, the prefix tree structure ensures fast insertion of new arrivals and deletion of scheduled requests, guaranteeing efficient scheduling during runtime. A formalized algorithm is demonstrated in Appendix A.2. For example, consider a system that can process two offline requests per batch with the following request queue: (What is ML, How to code, What is AI, How to debug). Under traditional FCFS scheduling, requests are processed in arrival order: (What is ML, How to code), (What is AI, How to debug), resulting in no prefix sharing opportunities. In contrast, PSM’s prefix-aware scheduling reorders requests as: (What is ML, What is AI), (How to code, How to debug), enabling KV-cache reuse through shared prefixes.
The PSM strategy demonstrates strong extendability. Under certain scenarios, the vanilla PSM strategy may lead to starvation for requests with minimal prefix-sharing potential, blocked by continuous arrivals for an extended period. To illustrate this challenge, consider a request queue: (What is ML, What is AI, How to code, What is DL). When new requests arrive with similar prefixes (What is LLM, What is DNN), a naive prefix-sharing policy would continuously prioritize requests sharing the What is prefix, potentially starving the How to code request indefinitely. This issue can be mitigated by an extended version of our PSM policy, combining maximum prefix sharing with request freshness by using a utility ratio to ensure a balance between efficiency and fairness. Based on the utility ratio, a new offline request would be selected from the DFS order of the prefix tree or the most stale request from a self-balanced binary search tree sorted by freshness. A detailed algorithm is in Appendix A.3. These enhancements can make the PSM strategy more practical for real-world deployments, retaining its efficiency while improving fairness and adaptability.
@@ 5 Performance Evaluation
Figure 4: HyGen improves serving throughput under varying SLOs.
@@ 5.1 Evaluation Setup
Implementation and Testbeds.
We implement HyGen on top of vLLM Kwon et al. (2023); vLLM Project (2023) and Sarathi-Serve Agrawal et al. (2024); Sarathi-Serve Project (2024), with 1,300 lines of additional code. We evaluate HyGen on two server configurations: one with 4 NVIDIA A100 GPUs (40GB VRAM each) and the other with 4 NVIDIA A40 GPUs (48GB VRAM each). Both servers have 64 CPU cores, 256GB DDR4 RAM, and a 1.5TB NVMe SSD.
Models and Workloads.
We evaluated HyGen using Llama2-7B Touvron et al. (2023) and Qwen-14B Bai et al. (2023) models on A100 and A40 GPUs, respectively.
For the online workload, we used the conversation trace from Azure LLM inference trace 2023 Patel et al. (2024), a one-hour production trace with real-world requests and timestamps. We randomly sampled the trace to achieve the desired QPS that suits our hardware serving capacity. Specifically, within a time duration of $T$ seconds, we would sample $T\times Q$ requests to suit a desired QPS $Q$. For the offline workload, we used arXiv summarization Cohan et al. (2018), a dataset for long document summarization.
@@
For interference evaluation, we focus on both Time to First Token (TTFT) and Time Between Tokens (TBT) and their respective mean and 99th percentile (P99) values. We measure throughput using tokens per second (TPS) and queries per second (QPS).
@@
For pure online inference, we use Sarathi Agrawal et al. (2024) as our baseline. For hybrid serving, we compare HyGen with two baselines for interference and throughput evaluation, respectively: (1) Sarathi++: We implement our online-first scheduling policy on Sarathi to support hybrid serving, including the request management and preemption handling policies introduced in Section 4.1. (2) HyGen*: To evaluate the throughput benefit of our HyGen design, we further improve Sarathi++ to an SLO-aware serving system, HyGen*. Besides inheriting the serving policies from Sarathi++, HyGen* serves offline requests at a specific offline QPS to control overall interference. The offline QPS is profiled using a similar design with the HyGen profiler to guarantee bounded SLO interference.
@@
Figure 6: Prefix Sharing Maximization policy achieves higher serving capacity.
Figure 7: SLO-aware profiler bridges the gap between batch latency and overall SLO.
@@
Figure 10: HyGen meets SLO under various online QPS settings.
Figure 11: HyGen is able to meet multiple SLOs simultaneously.
@@
Figure 13: Hygen meets SLOs and achieves higher throughput.
Figure 14: HyGen is robust to predictor accuracy.
Figure 15: Hygen dynamically adjusts offline throughput.
5.2 End-to-end Performance
HyGen respects latency requirements in co-locating requests. We tested HyGen on four SLO metrics, mean TBT, P99 TBT, mean TTFT, and P99 TTFT. For each metric, we set varying interference tolerance ratios as the SLO. Figure 3 shows that HyGen is able to control the interference and guarantees to meet specific SLOs across our settings. Compared with Sarathi++, an SLO-unaware system that yields the same result for all metrics and tolerance ratios, HyGen showcases efficient SLO-aware latency control.
HyGen improves serving throughput. Figure 4 shows the offline throughput of HyGen for various metrics and tolerance ratios. Through efficient request co-location, HyGen improves overall serving throughput by up to 3.87$\times$ compared to pure online serving. Under the same SLO, HyGen consistently achieves higher throughput compared to HyGen*, yielding up to 5.84×5.84$\times$5.84 × offline throughput gain. This verifies that our fine-grained latency predictor and SLO-aware profiler designs achieve higher serving efficiency compared to their simplified counterparts.
@@ 5.3 Performance Breakdown
We evaluate the accuracy of our latency predictor on Llama2-7B and Qwen-14B using Azure LLM Inference trace mixed with arXiv summarization dataset. Figure 5 shows that our latency predictor achieves a mean absolute percentage error of only 1.78% and 1.07%, confirming its high accuracy.
@@
To test HyGen’s compatibility with prefix sharing, we conducted a simulation experiment using Azure LLM Inference as the online trace and MMLU Hendrycks et al. (2021a; b) as the offline dataset on a Llama2-7B model. In our simulation, we deducted the shared prompt prefix length between consecutive offline requests to simulate prefix sharing. Figure 7 shows that HyGen yields up to 4×4$\times$4 × offline throughput gain with its prefix sharing maximization scheduling policy.
@@
To demonstrate the effect of HyGen’s SLO-aware profiler, we compared it with a simple strategy that sets the desired mean TBT SLO as the batch latency budget. Figure 7 shows the performance gap between individual batch latencies and overall mean TBT, illustrating how the SLO-aware profiler bridges this gap for controlled SLO in hybrid serving.
@@
Figure 8 shows a temporal throughput breakdown of HyGen. At runtime, HyGen dynamically adjusts offline throughput based on online workload and overall latency budget, batching offline requests more aggressively during online QPS troughs and reducing offline throughput during online bursts, harnessing compute resources in an adaptive manner.
@@ 5.4 Ablation Studies
To evaluate HyGen’s effectiveness in distributed inference, we deployed the Yi-34B model Young et al. (2024) on a server with 4 NVIDIA A40 GPUs using tensor-parallelism (TP) Shoeybi et al. (2019) and pipeline-parallelism (PP) Athlur et al. (2022); Narayanan et al. (2019); Huang et al. (2019) with degree 2 for each dimension. Using Azure LLM Inference and arXiv summarization workloads, Figure 9 shows that HyGen maintains its ability to meet SLOs and achieves higher offline throughput (up to 1.89×1.89$\times$1.89 ×) than the baseline.
@@
We further evaluate HyGen’s ability to meet stringent SLOs under varying online QPS settings on the four aforementioned metrics, each with 5% interference tolerance. Figure 10 shows that HyGen meets stringent SLOs for all metrics. We further demonstrate HyGen’s ability to meet multiple SLOs at the same time. By testing HyGen with a fixed P99 TTFT interference ratio (8%) and mean TBT interference ratios ranging from 10% to 50%, Figure 11 shows that at a lower mean TBT tolerance, HyGen’s performance is bounded by mean TBT SLOs; after reaching the fixed P99 TTFT SLO, mean TBT stops increasing in order to keep P99 TTFT under control.
@@
We further evaluate HyGen’s adaptability on two more experiments: The first using Mistral-7B model with Mooncake trace Qin et al. (2024), a trace containing request lengths and timestamps taken from real-world servers, as the online trace, and arXiv summarization as the offline trace; The second experiment uses Llama2-7B model with Azure LLM Inference trace and replaced the offline dataset with CNN/DailyMail summarization dataset See et al. (2017); Hermann et al. (2015). An analysis of the Mooncake trace is shown in Appendix B. Figure 12 and Figure 13 show that HyGen achieves superior performance than its counterparts under these settings.
@@
We tested HyGen’s robustness using several pre-trained LR latency predictors with varying prediction accuracy taken from other workloads and tested them on Azure LLM Inference trace and arXiv summarization dataset. Figure 15 shows how predictor accuracy (measured in mean absolute percentage error) affects offline throughput under the same P99 TBT SLO. HyGen remains robust across different accuracy settings. Additionally, our LR-based latency predictor is lightweight for training, with only $\sim$15ms training time for over 80,000 samples on CPUs.
Impact of online arrival rate.
Under a high online arrival rate, online requests may approach system serving capacity, limiting the headroom for request co-location. Figure 15 shows the effect of online QPS on offline throughput with 5% P99 TBT tolerance. As online load increases, HyGen adjusts offline throughput based on the system’s residual serving capacity while maintaining higher throughputs.
@@ 6 Conclusion
This paper introduces HyGen, an LLM serving system that enables efficient co-location of online and offline workloads. We employ control mechanisms to predict and manage interference impacts, and a scheduling policy to opportunistically schedule offline serving. Evaluation on production workloads demonstrates that HyGen improves serving throughput by 3.87-5.84$\times$ while maintaining strict latency SLOs.
@@ A.1 HyGen Two-phase Scheduling Algorithm
This section gives a detailed and formalized demonstration of the two-phase scheduling algorithm introduced in Section 4.1. In each scheduling step, the workflow invokes the SLO-aware scheduling process in Algorithm 1 twice (line 13 and line 18) to form a hybrid batch with online and offline request co-location while respecting latency and memory limits. To reduce the scheduling overhead of HyGen, we employ a message queue for asynchronous communication between the main process and the offline scheduler. After each scheduling step, the main process sends the metadata of batched requests to the message queue. The offline scheduler first calculates the expected status of each request based on scheduling decisions from the previous batch, and then runs a scheduling simulation to generate offline request scheduling decisions using our latency predictor and the profiled latency budget. The offline scheduling decisions are then sent back to the main process using the message queue and used for the next scheduling step. To support pipeline parallelization, a scheduling history archive of $K$ steps is kept by the offline scheduler for pipeline parallelization degree $K$, in order to have a holistic view of every request running in each pipeline stage at the time.
@@
3:  global $Q_{send}$, $Q_{recv}$ (message queues)